IDENTIFY and SET UP B=(1.35 x 10 ? T)(12i +3j - 4) is acting on a circular loop with area 4.45 cmThe loop has a current of 12Part B Find the components of initial magnetic torque of the loop Express your answers in newton-meters separated by commas.
Answer #1

7=(135x16 2 1) (ilî +3j -46). Area A=4.45 cm? = 4.45 x10-4 m2 Current I= 12.5A. Posptial magnetic moment. Ms-H(-0.81 +0.6j) PPartē Ho IA - 12.5x4.45 X10-4 = 5.562 X10-3 Am2. + H = -4 (-0.81 +061). 9 H =-5.562 x103(-0.82 +0.6). Hi = (4.45 Ĉ – 3.33 79tw, x = [4.45€ -1.33 7pU10 3 X 1:35x102(127+39-4R).. 77 =(.35810-5) (4.45 – 3.3377)(127 +3j -4R)] *7* = (1-35x10-5) [13.35î +